OpenStudy (jiteshmeghwal9):

If\[z_1=2+3i\]\[z_2=1+2i\]then \[z_1z_2-z_1^3=?\]

OpenStudy (empty):

Give it your best try and show your steps

OpenStudy (jiteshmeghwal9):

\[z_1.z_2=(2+3i)(1+2i)\]\[z_1z_2=(2-6)+i(4+3)\]\[z_1z_2=-4+7i\]now\[z_1^3=8+27i^3+54i^2+36i\]\[z_1^3=-46+27i^3+36i\]\[z_1z_2-z_1^3=42-29i+27i^3\]

OpenStudy (jiteshmeghwal9):

now wht next ?

ganeshie8 (ganeshie8):

Nice! \(i^3\) simplifies further right ?

OpenStudy (empty):

It looks almost right, except you seem to have messed up your terms here: \(z_1^3 = 8 + 12i + 18i^2 + 27i^3\) Which can also be simplified further with @ganeshie8 's comment. :D

OpenStudy (jiteshmeghwal9):

\(42-29i-27i\)=\(42-56i\)

ganeshie8 (ganeshie8):



OpenStudy (empty):

The picture is nice since all a cube power means is multiply something by itself three times, \(i^3 = i*i*i\) So simplifying this should be no big deal, since you know what \(i*i=-1\) already.

OpenStudy (jiteshmeghwal9):

the answer given in book is \(42-2i\). nt matching :/

ganeshie8 (ganeshie8):

must be an algebra error somewhere, just double check..

OpenStudy (jiteshmeghwal9):

yes\[z_1z_2-z_1^3=42-29i-27i^3\]\[=42-29i+27i=42-2i\]

OpenStudy (jiteshmeghwal9):

thanx

ganeshie8 (ganeshie8):

In case this is first time, it might be a bit exciting to notice that any power of \(i\) always simplifies to one of the numbers : \(\{\pm 1, ~\pm i\}\)

ganeshie8 (ganeshie8):

This holds : \[\large i^{n}\equiv i^{n\pmod{4}}\] In other words, subtracting/adding \(4\) from the exponent doesn't change the number